# Mutation-Guided Unit Test Generation with a Large Language Model

### v5

@@ I Introduction
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [naik2011software, fraser2014large, corno2004automatic, anand2013orchestrated]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [pacheco2007randoop, hamlet1994random, andrews2011genetic, oriat2005jartege, pacheco2007feedback] and search-based methods [fraser2011evosuite, rojas2016seeding, rojas2015combining, vivanti2013search]—to LLM-based approaches for generating unit tests [fan2023large, wang2024software, hou2024large].
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [chekam2017empirical, andrews2006using, inozemtseva2014coverage], high code coverage does not necessarily imply strong fault detection capability. Other research [chekam2017empirical, alshahwan2024automated, foster2025mutation] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation for LLM-generated test cases. For example, Foster et al. [foster2025mutation] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases from existing ones and generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [dakhel2024effective] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze why live mutants remain undetected, nor does it consider uncovered mutants, both of which are challenges we address in our research.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly maximizes the mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [tian2025fixing, wang2025projecttest, pan2024multi], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect use of assertion functions, including erroneous invocations and inappropriate parameter values.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Lastly, since achieving the maximum mutation score is inherently difficult, we introduce an iterative test-generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [yuan2024evaluating, gu2025llm, ryan2024code], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [just2014defects4j], it handles cross-module dependencies explicitly by simply providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage; instead, it focuses on improving mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [tian2025fixing, dakhel2024effective, siddiq2024using], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@
II Motivating Example and Informal Description
This section goes through a simple example to informally describe our approach, MutGen, which we will then precisely describe and formalize in the next section. The example Java code under test shown in Figure 1, which validates a date format, includes descriptive comments and input-output examples. For clarity and to save space, we retain only the most relevant parts of the comments in the figure. Most LLM-based techniques prompt LLMs using the following standard prompt:
“Generate a JUnit test for the following method to achieve high branch and line coverage.”
@@ II Motivating Example and Informal Description
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [foster2025mutation, chekam2017empirical, alshahwan2024automated], high coverage does not necessarily imply strong fault detection capability, for example, when measured as mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 1 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases such as February 29th on leap years, likely due to the absence of such examples in the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the comments provided within the program under test.
Further, when relying on mutation instead of code coverage, the generated inputs often fail to kill mutants effectively. For example, when the ConditionalsBoundary operator in PITest changes a conditional boundary and alters line 24 to:
@@
This change introduces a fault: the original code accepts day == 1 as valid, while the mutant incorrectly rejects it. We observe that LLM-generated test cases, when prompted with the original code, tend to focus on representative invalid inputs (e.g., day = 0) and valid values far from the boundary (such as day = 2 or higher). For instance, the model commonly produces a test such as:
@@
which fails in both the original function and mutant. However, it frequently fails to generate:
@@
To address these issues, we propose MutGen, an approach that improves the fault detection capability of LLM-generated test cases via a two-stage process: preprocessing and generation, which are detailed in Section III-B and Sections III-C-III-E, respectively.
II-A Preprocessing Stage
To address the first limitation above, MutGen removes comments from the target code and uses a prompt (shown in Figure 2) to instruct the LLM to summarize the method’s purpose and input format. The prompt used is as follows:
Figure 2: Prompt used for summarizing
“This Java function, ‘validDate‘, validates a given date string in the format ”mm-dd-yyyy” and returns ‘true‘ if the date is valid according to specific rules regarding month and day ranges, and ‘false‘ otherwise. The input date string must be non-empty and follow the exact specified format. If any of these conditions are not met, the function will return ‘false‘.”
To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with higher fault detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [pitest] to generate a mutation report, as shown in Figure 3. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 4) that guides the LLM in generating more effective test cases during the generation stage. For example, the mutation feedback extracted from the report for the mutant shown in Listing II is as follows:
21. conditional boundary at line 24, uncovered:
if (day<= 1||day>30) return false;
This means that the existing tests fail to cover the 21st mutant, which is generated by applying a conditional boundary change at line 24, as shown on the second line.
Figure 3: Example mutation report
II-B Generation Stage
Figure 4: Prompt used for generation
To achieve a high mutation score, MutGen employs a prompt augmented with mutation feedback, collected during the preprocessing stage, to guide Llama-3.3 in generating test cases, as illustrated below. For this HumanEval-Java subject, as shown in Figure 1, a single invocation of MutGen produces a test suite with 13 test cases, achieving a 70% mutation score across 30 mutants.
However, two of the generated test cases fail to execute due to runtime errors. To address this, MutGen incorporates a fixing mechanism that automatically repairs failing test cases, thereby improving both execution success rate and mutation score. In this example, MutGen successfully fixes one failing test case by invoking LLM with the prompt shown in Figure 5, resulting in a 13% increase in mutation score.
To further enhance effectiveness, MutGen introduces an iterative generation strategy. In each iteration, additional test cases are generated to target the remaining live and uncovered mutants that were not killed by the previous test suite. After two iterations, for the running example, MutGen achieves a 100% mutation score. During this process, LLMs may occasionally generate test cases with method names that conflict with existing method names. For example, as shown in Listing 1, the LLM generates a test case for the boundary value 04-01-2025 in April. Later, it may generate another test case for 04-31-2025, also intended to test boundary conditions. However, both may be assigned the same method name, leading to naming conflicts. To resolve this, MutGen incorporates an additional guideline, specifically addressing method name conflicts, into the prompt used for fixing (shown in Figure 5).
In contrast, as demonstrated in Section IV, for this subject in HumanEval-Java, using the standard prompt introduced at the beginning of this section, the LLM can only produce a test suite achieving a 53% mutation score, which remains unchanged even after four iterations.
Prompt Used for Fixing ## Failed Test The following test failed in the test suite: {{ failed_test with error message }} You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages. 1. Direct Invocation: • Always use assertion functions directly (e.g., assertEquals(value1, value2)) instead of prefixed calls like Assert.assertEquals(value1, value2). 2. Handling assertEquals Mismatches: • If an error follows the pattern: • test_file_name:line_number: expected:error_value but was:correct_value • And involves assertEquals, replace the expected value with the correct one. • Example: assertEquals(error_value, xxx); → assertEquals(correct_value, xxx); 3. Fixing assertTrue / assertFalse Errors: • If an error follows the pattern: • classname:line_number:null • And involves assertTrue or assertFalse, swap the function name while keeping all parameters unchanged. • Example: assertTrue(xxx); → assertFalse(xxx); 4. Handling ambiguous • If an error has the pattern: • reference to assertEquals is ambiguous • And involves assertEquals, add type casts to both parameters based on their respective types. • Example: assertEquals(expectedValue, actualValue); → assertEquals((long) expectedValue, (long) actualValue); 5. Method name conflicts • If an error follows the pattern: method is already defined in the class, • only append a number to the method name. • Example: public void test_null() → public void test_null1(); 6. Other errors • Try to fix the errors according to the error messages. ## Output Format in YAML The same applied to Figure 4.
Figure 5: Prompt used for fixing
III Our Approach
@@
In this section, we precisely describe and formalize our proposed approach, MutGen. As shown in Figure 6, MutGen takes a source program as input and returns test cases aimed at improving its fault-detection effectiveness by maximizing the test suite’s mutation score. We highlight all data flows in red in the figure. Steps 1, 2, and 5 are described in Sections III-B, III-C, and III-D, respectively. We do not include Section III-E as it reuses the components already shown in the figure. We begin with a high-level overview of MutGen, followed by detailed and formal descriptions of each individual component.
III-A Overview
@@
The preprocessing stage prepares information to guide test case generation, including code summarization and mutation feedback, as detailed in Section III-B. In the generation stage, the LLM is prompted with mutation feedback, such as descriptions of mutation operators and diff-like representations of mutants, to generate test cases.
@@
The final test suite consists of all passing test cases obtained by combining the filtered tests and the fixed tests. We denote this using the function ExeFilter, as follows:
@@
This process will be detailed in Section III-D.
@@
Our assumption, as reported in the mutation testing literature [chekam2017empirical, andrews2006using], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [foster2025mutation] and therefore enables the generation of more effective test suites. However, as revealed in recent work [gao2025prompt], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section III-E, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
III-B Preprocessing
@@
Code Summarization. As illustrated in Figure 6 and the example in Section II-A, the prompt instructs the LLM to summarize the purpose and the input format of the target code, resulting in a natural language description denoted as descdesc.
Mutation Feedback. Additionally, we apply a mutation testing tool to the code under test to generate mutation feedback. Since mutation testing tools rely on an existing test suite, we first create an empty test suite by mimicking a test function that returns directly. We then collect the mutation analysis results produced by the tool. As illustrated in Figure 6, the resulting mutation report includes information on live mutants, killed mutants, and uncovered mutants. Mutation feedback can be obtained by associating this information with the corresponding code.
As shown in the example in Section II-A, the corresponding mutation feedback entry is represented as:
@@
At the start of our approach, all mutants are initially alive, and the items of mutation feedback are formatted as shown in the above example.
III-C Prompting with Mutation Feedback
We define the prompt function with mutation feedback $\text{Prompt}_{\text{mut}}$ as:
$\text{Prompt}_{\text{mut}}$ $:\mathcal{R}\rightarrow\text{String}$
$\text{Prompt}_{\text{mut}}(\mathcal{R})$ $=\texttt{Format}\left(desc,{src,}\bigcup_{r_{i}\in\mathcal{R}}r_{i}\right)$

### v6

@@ I Introduction
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [49, 26, 16, 4]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [56, 33, 6, 53, 57] and search-based methods [25, 63, 62, 70]—to LLM-based approaches for generating unit tests [23, 72, 34].
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [12, 5, 36], high code coverage does not necessarily imply strong fault detection capability. Other research [12, 3, 24] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation for LLM-generated test cases. For example, Foster et al. [24] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases from existing ones and generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [17] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze why live mutants remain undetected, nor does it consider uncovered mutants, both of which are challenges we address in our research.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly maximizes the mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [68, 74, 58], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect use of assertion functions, including erroneous invocations and inappropriate parameter values.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Lastly, since achieving the maximum mutation score is inherently difficult, we introduce an iterative test-generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [78, 31, 64], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [40], it handles cross-module dependencies explicitly by simply providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage; instead, it focuses on improving mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [68, 17, 65], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@
II Methodology
This section presents our approach MutGen through a motivating example and a formal description. We first illustrate the limitations of standard LLM-based test generation using a running example. We then precisely describe and formalize our methodology, including the complete set of user prompts employed in our approach. To save space and focus on the essential context, we omit the system prompts and the standard formatting rules widely used in existing open-source tools [48] and test generation research [24, 14, 31], which are also available in our open-source implementation.
II-A Motivating Example and Intuition
This subsection goes through a simple example to illustrate the limitations of standard prompting and to motivate our approach. The example Java code under test shown in Figure 1, which validates a date format, includes descriptive comments and input-output examples. For clarity and to save space, we retain only the most relevant parts of the comments in the figure. Most LLM-based techniques prompt LLMs using the following standard prompt:
“Generate JUnit tests for the following method to achieve high branch and line coverage.”
1package original;
2class ValidDate {
3 /** You have to write a function ......
4 * The date is valid if all of the following rules are satisfied:
5 * 1-3. ......
6 * 4. The date should be in the format: $m$-$d$-yyyy
7 * for example:
8 * validDate("03-11-2000") => True
9 * validDate("06/04/2020") => False
10 * MORE EXAMPLES ...... **/
11 public static Boolean validDate(String date) {
12 if (date.length() != 10) {
13 return false;
14 }
15 String[] dateArr = date.split("-");
16 if (dateArr.length != 3) return false;
17 int month = Integer.parseInt(dateArr[0]);
18 int day = Integer.parseInt(dateArr[1]);
19 int year = Integer.parseInt(dateArr[2]);
20 if (month < 1 $|$ month > 12) { return false; }
21 if (month == 2) {
22 if (day < 1 $|$ day > 29) { return false; }
23 } else if (month == 4 $|$ month == 6 $|$ month == 9 $|$ month == 11) {
24 if (day < 1 $|$ day > 30) { return false; }
25 } else {
26 if (day < 1 $|$ day > 31) { return false; }
27 }
28 return true;
29 }
30}
@@ II-A Motivating Example and Intuition
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [24, 12, 3], high coverage does not necessarily imply strong fault-detection capability when measured by the mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 1 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases, such as February 29th in leap years, likely because such examples are absent from the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats, such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the program’s comments.
Furthermore, when relying on mutation rather than code coverage, the generated inputs often fail to effectively kill mutants. For example, when the ConditionalsBoundary mutation operator in PITest replaces < with <=, it changes the conditional boundary and modifies line 24 as follows:
@@
This change introduces a fault: the original code accepts day == 1 as valid, while the mutant incorrectly rejects it. We observe that LLM-generated test cases, when prompted with the original code, tend to focus on representative invalid inputs (e.g., day = 0) and on valid values far from the boundary (e.g., day = 2 or higher). For instance, the model commonly produces a test such as:
@@
which fails in both the original function and the mutant. However, it frequently fails to generate:
@@
To address these issues, we propose MutGen, an approach that improves the fault detection capability of LLM-generated test cases via a two-stage process: preprocessing and generation, which are detailed in Section II-C and Sections II-D-II-F, respectively.
II-B Formal Overview of MutGen
@@
The preprocessing stage prepares information to guide test case generation, including code summarization and mutation feedback, as detailed in Section II-C. In the generation stage, the LLM is prompted with mutation feedback, such as descriptions of mutation operators and diff-like representations of mutants, to generate test cases.
@@
The final test suite comprises all passing test cases obtained by combining the filtered and fixed tests. We denote this using the function ExeFilter, as follows:
@@
This process will be detailed in Section II-E.
@@
Our assumption, as reported in the mutation testing literature [12, 5], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [24] and therefore enables the generation of more effective test suites. However, as revealed in recent work [29], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section II-F, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
Figure 2 shows an overview of MutGen. MutGen takes a program under test as input and returns test cases aimed at improving its fault-detection effectiveness by maximizing the test suite’s mutation score. We highlight all data flows in red in the figure.
II-C Preprocessing Stage
@@
Code Summarization. To address the first limitation above, MutGen removes comments from the target code and uses a prompt (shown in Figure 3) to instruct the LLM to summarize the method’s purpose and input format. The prompt used is as follows:
Figure 3: Prompt used for summarizing
“This Java function, ‘validDate’, validates a given date string in the format “mm-dd-yyyy” and returns ‘true’ if the date is valid according to specific rules regarding month and day ranges, and ‘false’ otherwise. The input date string must be non-empty and follow the exact specified format. If any of these conditions are not met, the function will return ‘false’.”
Since our work focuses on improving the fault-detection capability of generated tests through mutation feedback, we directly use the summaries produced by the LLM without further optimization. Investigating how the size and quality of generated summaries affect test generation effectiveness is a promising direction for future work.
Mutation Feedback. To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with greater fault-detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [60] to generate a mutation report, as shown in Figure 4. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 5) that guides the LLM in generating more effective test cases during the generation stage.
For example, the mutation feedback extracted from the report for the mutant shown in Listing II-A is as follows:
21. conditional boundary at line 24, uncovered:
if (day<=1||day>30) return false;
This indicates that the existing tests do not cover the 21st mutant, generated by applying a conditional boundary change at line 24, specifically by replacing < with <=, as shown in the second line. Formally, this mutation feedback entry can be represented as:
@@
Mutations
12 1. negated conditional -> KILLED
13 1. replaced Boolean return with True for original/ValidDate::validDate -> KILLED
16 1. negated conditional -> KILLED
2. replaced Boolean return with True for original/ValidDate::validDate -> KILLED
20 1. negated conditional -> SURVIVED Covering tests
2. changed conditional boundary -> SURVIVED Covering tests
3. changed conditional boundary-> SURVIVED Covering tests
4. replaced Boolean return with True for original/ValidDate::validDate -> SURVIVED Covering tests
5. negated conditional-> SURVIVED Covering tests
21 1. negated conditional -> SURVIVED Covering tests
22 1. negated conditional -> SURVIVED Covering tests
2. changed conditional boundary -> SURVIVED Covering tests
3. changed conditional -> SURVIVED Covering tests
4. replaced Boolean return with True for original/ValidDate::validDate -> SURVIVED Covering tests
5. changed conditional boundary-> SURVIVED Covering tests
23 1. negated conditional -> SURVIVED Covering tests
2. negated conditional -> SURVIVED Covering tests
3. negated conditional -> SURVIVED Covering tests
4. negated conditional -> SURVIVED Covering tests
24 1. negated conditional -> NO_COVERAGE
2. changed conditional boundary -> NO_COVERAGE
3. changed conditional boundary -> NO_COVERAGE
4. replaced Boolean return with True for original/ValidDate::validDate -> NO_COVERAGE
5. negated conditional -> NO_COVERAGE
26 1. changed conditional boundary -> NO_COVERAGE
2. negated conditional -> NO_COVERAGE
3. changed conditional boundary -> NO_COVERAGE
4. replaced Boolean return with True for original/ValidDate::validDate -> NO_COVERAGE
5. negated conditional -> NO_COVERAGE
28 1. replaced Boolean return with False for original/ValidDate::validDate -> NO_COVERAGE
Figure 4: Example mutation report
II-D Prompting with Mutation Feedback
To maximize the mutation score, MutGen augments the prompt with mutation feedback collected during the preprocessing stage and uses it to guide Llama-3.3 in generating test cases. Figure 5 illustrates the resulting prompt for the HumanEval-Java subject shown in Figure 1. With a single invocation of MutGen, the generated test suite contains 13 test cases and achieves a mutation score of 70% over 30 mutants.
We define the prompt construction function as follows:
$\text{Prompt}_{\text{mut}}$ $:\mathcal{R}\rightarrow\text{String},$
$\text{Prompt}_{\text{mut}}(\mathcal{R})$ $=\texttt{Format}(desc,src,\bigcup_{r_{i}\in\mathcal{R}}r_{i}).$